OpenStudy (kamille):

Can anyone explain me how to do this? \[\int\limits_{0,5}^{1}\frac{ dx }{ x ^{2} }\]

OpenStudy (experimentx):

how do you do this\[ \int \frac{1}{x^2} dx\]

OpenStudy (kamille):

I know I need to insert numbers when I have done the integral, but I haven't learn at school how to integral when x is in denominator

OpenStudy (kamille):

thats why I am asking for help (for a rule?)

OpenStudy (anonymous):

Rewrite the integral as an exponential with x to a negative power, instead of as a fraction. Then use the basic power rule for integration.

OpenStudy (kamille):

well, once I have done it and my teacher told that I can't write like that, because it should be written with ln, Are you talking about this? \[\int\limits_{0,5}^{1}=x^{-2}dx\]

OpenStudy (experimentx):

when the power is -1 ... i.e \( \displaystyle \frac{1}{x} = x^{-1}\) that's a different case. when the power is not -1, use the usual power rule.

OpenStudy (kamille):

so I can write is as I have written above? : ) Thanks a lot, it always confuses me!

OpenStudy (anonymous):

Yes, what you've written is fine (just don't put the equal sign within the integral). The case with 1/x has an integral that gives you ln(x), but that is usually the only exception to the rule with fractions and x with negative exponents.

OpenStudy (kamille):

oh, okay! Thank you a lot, guys!